# Impact of existence and nonexistence of pivot on the coverage of empirical best linear prediction intervals for small areas

### v2

@@ Abstract
We advance the theory of parametric bootstrap in constructing highly efficient empirical best (EB) prediction intervals of small area means. The coverage error of such a prediction interval is of the order $O(m^{-3/2})$, where $m$ is the number of small areas to be pooled using a linear mixed normal model. In the context of an area level model where the random effects follow a non-normal known distribution except possibly for unknown hyperparameters, we analytically show that the order of coverage error of empirical best linear (EBL) prediction interval remains the same even if we relax the normality of the random effects by the existence of pivot for a suitably standardized random effects when hyperpameters are known. Recognizing the challenge of showing existence of a pivot, we develop a simple moment-based method to claim non-existence of pivot. We show that existing parametric bootstrap EBL prediction interval fails to achieve the desired order of the coverage error, i.e. $O(m^{-3/2})$, in absence of a pivot. We obtain a surprising result that the order $O(m^{-1})$ term is always positive under certain conditions indicating possible overcoverage of the existing parametric bootstrap EBL prediction interval. In general, we analytically show for the first time that the coverage problem can be corrected by adopting a suitably devised double parametric bootstrap. Our Monte Carlo simulations show that our proposed single bootstrap method performs reasonably well when compared to rival methods.
Keywords: Small area estimation, empirical Bayes, linear mixed model, best linear predictor
@@ 1 Introduction
The following two-level model, commonly referred to as the area-level model, has been extensively used in small area applications.
@@
Level 1 (Sampling model): $y_{i}\,|\,\theta_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{N}(\theta_{i},D_{i})$;
Level 2 (Linking model): $\theta_{i}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{G}(\bm{x_{i}^{\prime}\beta},A,\bm{\phi})$.
Here, $m$ represents the number of small areas, $\mathcal{G}$ is a fully parametric distribution, not necessarily normal, with mean $\text{E}(\theta_{i})=\bm{x_{i}^{\prime}\beta}$, variance $\text{Var}(\theta_{i})=A\geq 0$, and any additional parameters $\bm{\phi}$. Assuming the same normal distribution at level 1, Datta and Lahiri (1995) assumed that $\mathcal{G}$ is a scale mixture of normal distribution. Later, Bell and Huang (2006) and Xie et al. (2007) applied a $t$ distribution as a specific case of the scale mixture of normals at level 2, with mean $\bm{x_{i}^{\prime}\beta}$, variance $A$ and degrees of freedom $\phi$, to mitigate the influence of outliers. Fabrizi and Trivisano (2010) introduced two robust area-level models: the first assumes that $\theta_{i}$ follows an exponential power distribution with a mean $\bm{x_{i}^{\prime}\beta}$, variance $A$ and shape parameter $\phi$; the second model assumes a skewed exponential power distribution on level 2, with a mean $\bm{x_{i}^{\prime}\beta}$, variance $A$, shape parameter $\psi$ and the skewness parameter $\lambda$. Thus, in this case, the parameter vector $\bm{\phi}$ can be defined as $(\psi,\lambda)^{\prime}$.
In the above model, level 1 is used to account for the sampling distribution of the direct estimates $y_{i}$, which are weighted averages of observations from small area $i$. Level 2 links the true small area means $\theta_{i}$ to a vector of $p$ known auxiliary variables $\bm{x_{i}}=(x_{i1},\cdots,x_{ip})^{\prime}$, often obtained from various administrative records. The parameters $\bm{\beta}$ and $A$ of the linking model are generally unknown and are estimated from the available data. As in other papers on the area-level model, the sampling variances $D_{i}$ are assumed to be known. In practice, $D_{i}$’s are estimated using a smoothing technique such as the ones given in Fay and Herriot (1979), Otto and Bell (1995), Ha et al. (2014) and Hawala and Lahiri (2018).
@@
$y_{i}=\theta_{i}+e_{i}=\bm{x_{i}^{\prime}\beta}+u_{i}+e_{i},\;i=1,\cdots,m,$ (1.1)
@@
$\tilde{\theta_{i}}^{\rm{BLP}}=(1-B_{i})y_{i}+B_{i}\bm{x_{i}^{\prime}\beta},$ (1.2)
where $B_{i}=D_{i}/(A+D_{i})$, and the BLP minimizes the mean squared prediction error (MSPE) among all linear predictors of $\theta_{i}$. The variance of the prediction error $(\theta_{i}-\tilde{\theta_{i}}^{\rm{BLP}})$ is denoted by $g_{1i}(A)\coloneq\text{Var}(\theta_{i}-\tilde{\theta_{i}}^{\rm{BLP}})=AD_{i}/(A+D_{i})$. If $A$ is known, one can obtain the standard weighted least squares estimator of $\bm{\beta}$, denoting as $\tilde{\bm{\beta}}(A)$. Replacing $\bm{\beta}$ in (1.2) by $\tilde{\bm{\beta}}=\tilde{\bm{\beta}}(A)$, a best linear unbiased prediction (BLUP) estimator of $\theta_{i}$ is given by
$\tilde{\theta_{i}}^{\text{BLUP}}=(1-B_{i})y_{i}+B_{i}\bm{x_{i}^{\prime}\tilde{\bm{\beta}}},$ (1.3)
@@
$\hat{\theta}_{i}^{\rm{EBLUP}}=(1-\hat{B}_{i})y_{i}+\hat{B}_{i}\bm{x_{i}^{\prime}\hat{\beta}},$ (1.4)
where $\hat{B}_{i}=D_{i}/(\hat{A}+D_{i})$ and $\hat{\bm{\beta}}=\tilde{\bm{\beta}}(\hat{A})$, and $\hat{A}$ is a consistent estimator of $A$ for large $m$. After plugging in the estimator $\hat{A}$, we have $g_{1i}(\hat{A})=\hat{A}D_{i}/(\hat{A}+D_{i})$. To simplify the notation, throughout the remainder of this paper, we will use $g_{1i}$ and $\hat{g}_{1i}$ to denote $g_{1i}(A)$ and $g_{1i}(\hat{A})$, respectively. Point prediction using EBLUP and the associated mean squared prediction error (MSPE) estimation have been studied extensively. See Rao and Molina (2015) and Jiang and Nguyen (2007) for a detailed discussion on this subject.
In this paper, we consider prediction interval approximates of small area means $\theta_{i}$. A prediction interval of $\theta_{i}$, denoted by $I_{i}$, is called a $100(1-\alpha)\%$ interval for $\theta_{i}$ if $\mathrm{P}(\theta_{i}\in I_{i}|\bm{\beta},A,\bm{\phi})=1-\alpha$, for any fixed $\bm{\beta}\in\mathcal{R}^{p}$, $A\in\mathcal{R}^{+}$ and $\bm{\phi}\in\Theta_{\bm{\phi}}$, the parameter space of $\bm{\phi}$. The probability $\mathrm{P}$ is with respect to the area-level model. There are several options for constructing interval estimates for $\theta_{i}$. The prediction interval based on only the Level 1 model for the observed data is given by $y_{i}\pm z_{\alpha/2}\sqrt{D_{i}}$, where $z_{\alpha/2}$ is the $100(1-\alpha/2)$th standard normal percentile. While the coverage probability for this direct interval is $1-\alpha$, it is not efficient when $D_{i}$ is large as in the case of small area estimation. Hall and Maiti (2006) considered the interval based the regression synthetic estimator: $(\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{A}},\bm{x_{i}^{\prime}}\hat{\bm{\beta}}+b_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{A}})$, where $b_{\hat{\alpha}_{l}^{s}}$ and $b_{\hat{\alpha}_{u}^{s}}$ are obtained using a parametric bootstrap method (described in detail in a later section). However, this interval is synthetic in the sense that it is constructed using synthetic regression estimator of $\theta_{i}$ and its associated uncertainty measure, which are not area specific in the outcome variable and hence may cause larger length of the confidence interval.
@@
A function of the data and parameters is called to be a pivot if its distribution does not depend on any unknown quantity (Shao, 2008; Hall, 2013). When we have a linear mixed normal model on (1.1) , $(\theta_{i}-\tilde{\theta}_{i}^{\rm{BLP}})/\sqrt{g_{1i}}$ is a standard normal pivot. The traditional method of interval estimation for $\theta_{i}$ is of the form $\rm{EBLUP}\pm z_{\alpha/2}\sqrt{\text{mspe}}$, where mspe is an estimate of the true mean squared prediction error (MSPE) of the EBLUP. Unfortunately, $(\theta_{i}-\hat{\theta}_{i}^{\text{EBLUP}})/\sqrt{\hat{g}_{1i}}$ is not a pivot and this traditional approach produces too short or too long intervals. The coverage error of such interval is of the order $O(m^{-1})$, not accurate enough in most small area applications. Recognizing that $(\theta_{i}-\hat{\theta}_{i}^{\text{EBLUP}})/\sqrt{\hat{g}_{1i}}$ does not follow a standard normal distribution, Chatterjee et al. (2008) and Li and Lahiri (2010) developed a parametric bootstrap method to approximate its distribution and obtained a EB prediction interval for $\theta_{i}$ in linear mixed normal models. They showed such interval has the coverage error of the order $O(m^{-3/2})$. However, the property remains unknown for non-normal distributed random effects.
In this paper, one main aim is to bring out the virtues of pivoting or rescaling, which can decrease the dependence of our statistics on unknown parameters and yield improved prediction interval approximation for small areas under the general model (1.1). Analogous to Chatterjee et al. (2008), we propose parametric bootstrap methods to approximate the distribution of a suitably centered and scaled EBLUP under the general model (1.1), and apply it to construct a $100(1-\alpha)\%$ prediction interval estimation of $\theta_{i}$. Here, we define an interval based on the EBLUP of $\theta_{i}$ as an Empirical Best Linear (EBL) prediction interval. Specifically, we introduce these two key quantities: one is the centered and scaled EBLUP following the $F_{1i}$ distribution, which can be expressed as below
@@
In Section 2, we introduce a list of notation and regularity conditions used throughout the paper. In Section 3, we propose a single parametric bootstrap EBL prediction interval for small area means in the context of an area-level model, where the random effects follow a general (non-normal) known distribution, except possibly for unknown hyperparameters. We analytically demonstrate that the coverage error order of the EBL prediction interval remains the same as in Chatterjee et al. (2008), even when relaxing the normality of random effects through the existence of a pivot for suitably standardized random effects when hyperparameters are known. However, when a pivot does not exist, the EBL prediction interval fails to achieve the desired coverage error order, i.e. $O(m^{-3/2})$. Surprisingly, we find that the order $O(m^{-1})$ term is always positive under certain conditions, indicating potential overcoverage in the current single parametric bootstrap EBL prediction interval. Recognizing the challenge of showing existence of a pivot, we develop a simple moment-based method to claim non-existence of pivot. In Section 4, we propose a double parametric bootstrap method under a general area level model and, for the first time, analytically show that this approach can correct the coverage problem. In Section 5, we compare our proposed EBL prediction interval methods with the direct method, other traditional approaches, and the parametric bootstrap prediction interval proposed by Hall and Maiti (2006).
@@ 2 A list of notations and regularity conditions
$\bm{\tilde{\beta}}=(\bm{X^{\prime}\bm{\Sigma}^{-1}X)^{-1}X^{\prime}\bm{\Sigma}^{-1}Y}$, weighted least square estimator of $\bm{\beta}$ with known $A$;
$\bm{P=\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}X(X^{\prime}\bm{\Sigma}^{-1}X)^{-1}X^{\prime}\bm{\Sigma}^{-1}}$;
@@
    $\text{sup}_{i\geq 1}h_{ii}\equiv\bm{x_{i}^{\prime}(X^{\prime}X)^{-1}x_{i}}=O(m^{-1})$;
@@
    $\text{E}|u_{i}|^{8+\delta}<\infty$ for $10<\delta<$ 1;
@@
    $\begin{split}\operatorname{E}[(\bm{\hat{\nu}}-\bm{\nu})^{\prime}\partial F_{2i}(x;\bm{\nu})/\partial\bm{\hat{\nu}}]=&\;O(m^{-1}),\\ \operatorname{E}\left[(\bm{\hat{\nu}}-\bm{\nu})^{\prime}\Bigl{\{}\partial^{2}F_{2i}(x;\bm{\nu})/\partial\bm{\hat{\nu}}\partial\bm{\hat{\nu}^{\prime}}\Bigr{\}}(\bm{\hat{\nu}}-\bm{\nu})\right]=&\;O(m^{-1}),\\ \text{E}\lVert(\bm{\hat{\nu}}-\bm{\nu})\rVert^{3}=\sum_{i,j,k}\text{E}[(\hat{\nu}_{i}-\nu_{i})(\hat{\nu}_{j}-\nu_{j})(\hat{\nu}_{k}-\nu_{k})]=&\;o(m^{-1}),\end{split}$
@@ 3 Single parametric bootstrap
    Conditionally on the data $(\bm{X,Y})$, draw $\theta_{i}^{*}$ for $i=1,\cdots,m$, independently from the distribution $\mathcal{G}(\bm{x_{i}}^{\prime}\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})$;
@@
    $\begin{split}\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})&=1-\alpha/2\\ \mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{l}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})&=\alpha/2,\end{split}$ (3.1)
@@
One of our main results from the algorithm above is that we relax the normality of the random effects by the existence of a pivot $H_{i}(\bm{\beta},A)$, the prediction interval $\hat{I}_{i}^{s}$ obtained above still has a high degree of coverage accuracy. That is, it brings the coverage error down to $O(m^{-3/2})$.
Theorem 3.1.
Under regularity conditions, for a preassigned $\alpha\in(0,1)$ and arbitrary $i=1,\cdots,m$, when $H_{i}(\bm{\beta},A)$ is a pivot, we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=\mathrm{P}\left(\hat{\theta}_{i}+q_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{g}_{1i}}\leq\theta_{i}\leq\hat{\theta}_{i}+q_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{g}_{1i}}\right)=1-\alpha+O(m^{-3/2}),$ (3.2)
@@
The proof of Theorem 3.1 is given in the Appendix A. An example of Theorem 3.1 is that when $\mathcal{G}(\bm{x_{i}^{\prime}\beta},A,\bm{\phi})$ is a normal distribution, $\mathcal{N}(\bm{x_{i}^{\prime}\beta},A)$, using Theorem 3.2 of Chatterjee et al. (2008), we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+O(m^{-3/2}).$ (3.3)
Proposition 3.1.
When $H_{i}(\bm{\beta},A)$ is not a pivot, we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+O(m^{-1}),$ (3.4)
where $\hat{I}^{s}_{i}$ is obtained from the single parametric bootstrap procedure described above.
The proof is given in the Appendix B.
Proposition 3.2.
@@
[…]
    This condition is satisfied for some continuous distributions. For instance, when $u_{i}$ follows a logistic or $t$ distribution with known degrees of freedom. The only unknown parameter of $F_{2i}(x;\bm{\nu})$ is the variance $A$. As Remark 1 indicates that the kurtosis of $H_{i}(\bm{\beta},A)$ is a decreasing function of $A$, it is not difficult to show that $\partial F_{2i}(x;A)/\partial A>0$ and $\partial^{2}F_{2i}(x;A)/\partial A^{2}<0$;
@@
Under these conditions, the prediction interval (3.4) has an overcoverage property. More specifically, we can rewrite (3.4) as below:
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+T_{1}+o(m^{-1}),$ (3.5)
@@
See detailed proof in the Appendix C. The proposition indicates that under the regularity conditions, the prediction intervals conducted by the proposed single parametric bootstrap can produce higher coverage than the nominal coverage with a non pivot $H_{i}(\bm{\beta},A)$ up to the order of $O(m^{-1})$, which could be beneficial for practitioners without considering other properties of prediction intervals.
Remark 1.

### v3

@@ Abstract
We advance parametric bootstrap theory to construct highly efficient empirical best prediction intervals for small area means, achieving a coverage error rate of $O(m^{-3/2})$, where $m$ is the number of areas modeled by a linear mixed normal model. For a general mixed effect model with random effects from a known but non-normal distribution (with unknown hyperparameters), we show analytically that the empirical best linear (EBL) prediction interval maintains the same coverage error order, assuming the existence of a pivot for standardized random effects with known hyperparameters. Recognizing challenges in proving pivot existence, we develop a simple moment-based method to claim non-existence of pivot. We find that without a pivot, the parametric bootstrap EBL prediction interval fails to reach the desired $O(m^{-3/2})$ coverage error. We obtain a surprising result that the order $O(m^{-1})$ term is always positive under certain conditions indicating possible overcoverage of the existing parametric bootstrap EBL prediction interval. In general, we analytically show for the first time that the coverage problem can be corrected by adopting a suitably devised double parametric bootstrap. Our Monte Carlo simulations show that our proposed single bootstrap method performs reasonably well when compared to rival methods.
Keywords: Small area estimation; empirical Bayes; linear mixed model; best linear predictor.
@@ 1 Introduction
Small area estimation plays a vital role in survey applications, where public and private agencies depend on accurate statistical inference. While previous research has extensively addressed point prediction and associated mean squared prediction error (MSPE), interval estimation has typically been limited to specific cases, such as the linear mixed normal model. In this paper, we propose parametric bootstrap methods for constructing empirical best prediction intervals under a general area-level model. We analytically demonstrate that the proposed methods achieve high efficiency in terms of coverage error, apply to a much broader class of models than existing approaches, and avoid the need for complex analytical derivations.
The following two level model, commonly referred to as the area level model, has been extensively used in small area applications.
@@
Level 1 (Sampling model): $y_{i}\,|\,\theta_{i}\overset{\mathrm{ind}}{\sim}\mathcal{N}(\theta_{i},D_{i})$;
Level 2 (Linking model): $\theta_{i}\overset{\mathrm{ind}}{\sim}\mathcal{G}(\bm{x_{i}^{\prime}\beta},A,\bm{\phi})$.
Here, $m$ represents the number of small areas. The level 1 model is used to account for the sampling distribution of the direct estimates $y_{i}$, which are weighted or unweighted averages of observations from small area $i$. The level 2 model links the true small area means $\theta_{i}$ to a vector of $p$ known auxiliary variables $\bm{x_{i}}=(x_{i1},\cdots,x_{ip})^{\prime}$, often obtained from various administrative records. The coefficient vector $\bm{\beta}\in\mathcal{R}^{p}$ and model variance parameter $A$ of the linking model are generally unknown and are estimated from the available data. As in other papers on the area-level model, the sampling variances $D_{i}$ are assumed to be known. In practice, $D_{i}$’s are estimated using a smoothing technique such as the ones given in Fay and Herriot (1979), Otto and Bell (1995), Ha et al. (2014) and Hawala and Lahiri (2018).
The classical area-level model introduced by Fay and Herriot (1979) assumes normality at both levels. The normality assumption at Level 1 is often justified by the Central Limit Theorem, as $y_{i}$ typically represents a sample summary (Rao and Molina, 2015; Jiang and Torabi, 2022). However, the normality assumption at Level 2 is more challenging to justify. Therefore, in this work, we consider that $\mathcal{G}$ is a fully parametric distribution, not necessarily normal, with mean $\text{E}(\theta_{i})=\bm{x_{i}^{\prime}\beta}$, variance $\text{Var}(\theta_{i})=A\geq 0$, and any additional parameters $\bm{\phi}$. For instance, Datta and Lahiri (1995) assumed that $\mathcal{G}$ is a scale mixture of normal distribution. Later, Bell and Huang (2006) and Xie et al. (2007) applied a $t$ distribution as a specific case of the scale mixture of normals at level 2, with mean $\bm{x_{i}^{\prime}\beta}$, variance $A$ and degrees of freedom $\phi$, to mitigate the influence of outliers. Fabrizi and Trivisano (2010) introduced two robust area-level models: the first assumes that $\theta_{i}$ follows an exponential power distribution with a mean $\bm{x_{i}^{\prime}\beta}$, variance $A$ and shape parameter $\phi$; the second model assumes a skewed exponential power distribution on level 2, with a mean $\bm{x_{i}^{\prime}\beta}$, variance $A$, shape parameter $\psi$ and the skewness parameter $\lambda$. Thus, in this case, the parameter vector $\bm{\phi}$ can be defined as $(\psi,\lambda)^{\prime}$. Similarly, Jiang and Torabi (2022) considered that $\theta_{i}$ follows a skewed normal distribution with mean $\bm{x_{i}^{\prime}\beta}$, variance $A$, and skewness parameter $\phi$.
@@
$y_{i}=\theta_{i}+e_{i}=\bm{x_{i}^{\prime}\beta}+u_{i}+e_{i},\;i=1,\cdots,m,$ (1)
@@
$\tilde{\theta_{i}}^{\rm{BLP}}=(1-B_{i})y_{i}+B_{i}\bm{x_{i}^{\prime}\beta},$ (2)
where $B_{i}=D_{i}/(A+D_{i})$, and the BLP minimizes the mean squared prediction error (MSPE) among all linear predictors of $\theta_{i}$. Note that having a closed-form expression not only facilitates theoretical analysis but also simplifies practical implementation. The variance of the prediction error $(\theta_{i}-\tilde{\theta_{i}}^{\rm{BLP}})$ is denoted by $g_{1i}(A)\coloneq\text{Var}(\theta_{i}-\tilde{\theta_{i}}^{\rm{BLP}})=AD_{i}/(A+D_{i})$. If $A$ is known, one can obtain the standard weighted least squares estimator of $\bm{\beta}$, denoting as $\tilde{\bm{\beta}}(A)$. Replacing $\bm{\beta}$ in (2) by $\tilde{\bm{\beta}}=\tilde{\bm{\beta}}(A)$, a best linear unbiased prediction (BLUP) estimator of $\theta_{i}$ is given by
$\tilde{\theta_{i}}^{\text{BLUP}}=(1-B_{i})y_{i}+B_{i}\bm{x}_{i}^{\prime}\tilde{\bm{\beta}}$, (3)
@@
$\hat{\theta}_{i}^{\rm{EBLUP}}=(1-\hat{B}_{i})y_{i}+\hat{B}_{i}\bm{x}_{i}^{\prime}\hat{\bm{\beta}}$, (4)
[…]
@@
A function of the data and parameters is called to be a pivot if its distribution does not depend on any unknown quantity (Shao, 2008; Hall, 2013). When we have a linear mixed normal model on (1) , $(\theta_{i}-\tilde{\theta}_{i}^{\rm{BLP}})/\sqrt{g_{1i}}$ is a standard normal pivot. The traditional method of interval estimation for $\theta_{i}$ is of the form $\rm{EBLUP}\pm z_{\alpha/2}\sqrt{\text{mspe}}$, where mspe is an estimate of the true mean squared prediction error (MSPE) of the EBLUP. Unfortunately, $(\theta_{i}-\hat{\theta}_{i}^{\text{EBLUP}})/\sqrt{\hat{g}_{1i}}$ is not a pivot and this traditional approach produces too short or too long intervals. The coverage error of such interval is of the order $O(m^{-1})$, not accurate enough in most small area applications. Recognizing that $(\theta_{i}-\hat{\theta}_{i}^{\text{EBLUP}})/\sqrt{\hat{g}_{1i}}$ does not follow a standard normal distribution, Chatterjee et al. (2008) and Li and Lahiri (2010) developed a parametric bootstrap method to approximate its distribution and obtained a EB prediction interval for $\theta_{i}$ in linear mixed normal models. They showed such interval has the coverage error of the order $O(m^{-3/2})$. However, the property remains unknown for non-normal distributed random effects.
In this paper, one main aim is to bring out the virtues of pivoting or rescaling, which can decrease the dependence of our statistics on unknown parameters and yield improved prediction interval approximation for small areas under the general model (1). Analogous to Chatterjee et al. (2008), we propose parametric bootstrap methods to approximate the distribution of a suitably centered and scaled EBLUP under the general model (1), and apply it to construct a $100(1-\alpha)\%$ prediction interval estimation of $\theta_{i}$. In fact, our theory does not require normality at either level of the area level model.
Here, we define an interval based on the EBLUP of $\theta_{i}$ as an Empirical Best Linear (EBL) prediction interval. Specifically, we introduce these two key quantities: one is the centered and scaled EBLUP following the $F_{1i}$ distribution, which can be expressed as below:
@@
In Section 2, we introduce a list of notation and regularity conditions used throughout the paper. In Section 3, we propose a single parametric bootstrap EBL prediction interval for small area means in the context of an area-level model, where the random effects follow a general (non-normal) known distribution, except possibly for unknown hyperparameters. We analytically demonstrate that the coverage error order of the EBL prediction interval remains the same as in Chatterjee et al. (2008), even when relaxing the normality of random effects through the existence of a pivot for suitably standardized random effects when hyperparameters are known. However, when a pivot does not exist, the EBL prediction interval fails to achieve the desired coverage error order, i.e. $O(m^{-3/2})$. Surprisingly, we find that the order $O(m^{-1})$ term is always positive under certain conditions, indicating potential overcoverage in the current single parametric bootstrap EBL prediction interval. Recognizing the challenge of showing existence of a pivot, we develop a simple moment-based method to claim non-existence of pivot. In Section 4, we propose a double parametric bootstrap method under a general area level model and, for the first time, analytically show that this approach can correct the coverage problem. In Section 5, we compare our proposed EBL prediction interval methods with the direct method, other traditional approaches, and the parametric bootstrap prediction interval proposed by Hall and Maiti (2006). In Section 6, we demonstrate our proposed methods using real life data from the 1989 Small Area Income and Poverty Estimates (SAIPE) program.
@@ 2 A list of notations and regularity conditions
$\tilde{\bm{\beta}}=(\bm{X}^{\prime}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{\prime}\bm{\Sigma}^{-1}\bm{Y}$, weighted least square estimator of $\bm{\beta}$ with known $A$;
$\bm{P}=\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\bm{X}(\bm{X}^{\prime}\bm{\Sigma}^{-1}\bm{X})^{-1}\bm{X}^{\prime}\bm{\Sigma}^{-1}$;
@@
    $\text{sup}_{i\geq 1}h_{ii}\equiv\bm{x}_{i}^{\prime}(\bm{X}^{\prime}\bm{X})^{-1}\bm{x}_{i}=O(m^{-1})$;
@@
    $\text{E}|u_{i}|^{8+\delta}<\infty$ for $0<\delta<1$;
@@
    $\text{E}[(\hat{\bm{\nu}}-\bm{\nu})^{\prime}\partial F_{2i}(x;\bm{\nu})/\partial\hat{\bm{\nu}}$] $=$ $O(m^{-1}),$
    $\text{E}\left[(\hat{\bm{\nu}}-\bm{\nu})^{\prime}\Bigl\{\partial^{2}F_{2i}(x;\bm{\nu})/\partial\hat{\bm{\nu}}\partial\hat{\bm{\nu}^{\prime}}\Bigr\}(\hat{\bm{\nu}}-\bm{\nu})\right]$ $=$ $O(m^{-1}),$
    $\text{E}\lVert(\hat{\bm{\nu}}-\bm{\nu})\rVert^{3}=\sum_{i,j,k}\text{E}[(\hat{\nu}_{i}-\nu_{i})(\hat{\nu}_{j}-\nu_{j})(\hat{\nu}_{k}-\nu_{k})]$ $=$ $o(m^{-1}),$
@@ 3 Single parametric bootstrap
[…]
@@
    $\begin{split}\mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{u}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})&=1-\alpha/2\\ \mathrm{P}^{*}(H_{i}(\hat{\bm{\beta}}^{*},\hat{A}^{*})\leq q_{\alpha_{l}}\,|\,\hat{\bm{\beta}},\hat{A},\hat{\bm{\phi}})&=\alpha/2,\end{split}$ (5)
@@
One of our main results from the algorithm above is that when we relax the normality of the random effects by the existence of a pivot $H_{i}(\bm{\beta},A)$, the prediction interval $\hat{I}_{i}^{s}$ obtained above still has a high degree of coverage accuracy. That is, it brings the coverage error down to $O(m^{-3/2})$.
Theorem 1
Under regularity conditions R1-R5, for a preassigned $\alpha\in(0,1)$ and arbitrary $i=1,\cdots,m$, when $H_{i}(\bm{\beta},A)$ is a pivot, we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=\mathrm{P}\left(\hat{\theta}_{i}+q_{\hat{\alpha}_{l}^{s}}\sqrt{\hat{g}_{1i}}\leq\theta_{i}\leq\hat{\theta}_{i}+q_{\hat{\alpha}_{u}^{s}}\sqrt{\hat{g}_{1i}}\right)=1-\alpha+O(m^{-3/2}),$ (6)
@@
The proof of Theorem 1 is given in the online Appendix A. A special example of Theorem 1 is that when $\mathcal{G}(\bm{x_{i}^{\prime}\beta},A,\bm{\phi})$ is a normal distribution, $\mathcal{N}(\bm{x_{i}^{\prime}\beta},A)$, using Theorem 3.2 of Chatterjee et al. (2008), we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+O(m^{-3/2}).$ (7)
Remark 1
We illustrate the existence of a pivot for the $H_{i}(\bm{\beta},A)$ in a setting where the $\theta_{i}$ at level 2 follows a non-normal distribution. Consider the following hierarchical model, which generalizes the model in Ghosh and Lahiri (1987) by incorporating covariates in the mean structure:
$\text{Level 1}:$ $\;y_{i}|\theta_{i},h\overset{\mathrm{ind}}{\sim}\mathcal{N}(\theta_{i},h^{-1});$
$\text{Level 2}:$ $\;\theta_{i}|h\overset{\mathrm{ind}}{\sim}\mathcal{N}(\bm{x}_{i}^{\prime}\bm{\beta},(\eta h)^{-1}),$
where $h$ has a proper prior distribution. Assume $h\sim\Gamma(\alpha,\lambda)$ with known shape parameter $\alpha>1$ and known scale parameter $\lambda$. In this case, the marginal distribution of $\theta_{i}$ is a $t$ distribution with $2\alpha$ degrees of freedom, mean $\bm{x}_{i}^{\prime}\bm{\beta}$, and scale parameter $\sqrt{\lambda/(\eta\alpha)}$: $\theta_{i}\sim t_{2\alpha}(\bm{x}_{i}^{\prime}\bm{\beta},\sqrt{\lambda/(\eta\alpha)})$. Moreover, the $H_{i}(\bm{\beta},A)$ also follows a $t$ distribution with known parameters:
$H_{i}(\beta,A)\sim t_{2\alpha}(0,\lambda\sqrt{(\alpha-1)/\alpha}).$
Since the distribution of $H_{i}(\bm{\beta},A)$ does not depend on unknown parameters, it is a pivot.
Proposition 1
Given a non-pivot $H_{i}(\bm{\beta},A)$, and under regularity conditions R1-R6, we have
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+O(m^{-1}),$ (8)
where $\hat{I}_{i}^{s}$ is obtained from the single parametric bootstrap procedure described above.
The proof is given in the online Appendix B.
Proposition 2
@@
    $\partial F_{2i}(x;\bm{\nu})/\partial\nu_{i}>0,\;1\leq i\leq s$ and $\lambda_{\rm{max}}(\partial^{2}F_{2i}(x;\bm{\nu})/\partial\bm{\nu}\partial\bm{\nu}^{\prime})<0$, where $\lambda_{\rm{max}}$ means the largest eigenvalue.
    This condition is satisfied for some continuous distributions. For instance, when $u_{i}$ follows a logistic or $t$ distribution with known degrees of freedom. The only unknown parameter of $F_{2i}(x;\bm{\nu})$ is the variance $A$. As indicated in Remark 2, the kurtosis of $H_{i}(\bm{\beta},A)$ is a decreasing function of $A$, and thus it is not difficult to show that $\partial F_{2i}(x;A)/\partial A>0$ and $\partial^{2}F_{2i}(x;A)/\partial A^{2}<0$;
@@
Under these conditions with the regularity conditions R1-R6 , the prediction interval (8) exhibits an overcoverage property. More specifically, we can rewrite (8) as below:
$\mathrm{P}(\theta_{i}\in\hat{I}_{i}^{s})=1-\alpha+T_{1}+o(m^{-1}),$ (9)
@@
See detailed proof in the online Appendix C. The proposition indicates that under the regularity conditions, the prediction intervals conducted by the proposed single parametric bootstrap can produce higher coverage than the nominal coverage with a non pivot $H_{i}(\bm{\beta},A)$ up to the order of $O(m^{-1})$, which could be beneficial for practitioners without considering other properties of prediction intervals.
Remark 2